(1 point) We consider the initial value problem 4xy + 4xy +9y = 0, y1) = 1, y(1) = -3 By looking for solutions in the form
Answer #1

- Now use y(12=) lac, lop 13 en (1)) + 2 sin (3ena) 1=ci log (o)tee sinco) Tol;l to => 10,=1 Again use y (1)=-3 -3=-3 csin (